OpenStudy (anonymous):

cheyanne plans to put $150 into a savings account. she can place her money into an account represented by f(x)=5x+150, or into another account represented by g(x)=150(1.03)^x. which account has the highest value in 3 years? which account has the highest value in 10 years?

OpenStudy (anonymous):

these are the answer choices: 1. f(x) has the highest value in 3 years; g(x) has the highest value in 10 years 2. g(x) has the highest value in 3 years; f(x) has the highest value in 10 years 3. f(x) has the highest value in 3 years; f(x) has the highest value in 10 years 4.g(x) has the highest value in 3 years; g(x) has the highest value in 10 years

Vocaloid (vocaloid):

x is the number of years, right? so first, find f(x) when x = 3 and g(x) when x = 3

OpenStudy (anonymous):

Alright I would plug those stuff into calculator

OpenStudy (anonymous):

Set the x to be 3 years and 10 years and see if either of them turns out greater than the other.

OpenStudy (anonymous):

yea i got f(x)= 165 g(x)= 3,687

Vocaloid (vocaloid):

check g(x) again

Vocaloid (vocaloid):

g(x)=150(1.03)^x = 150(1.03)^3 = ?

OpenStudy (anonymous):

163?

Vocaloid (vocaloid):

right, so which one is higher?

OpenStudy (anonymous):

g(x)

OpenStudy (anonymous):

For 3 years f(x)=5x+150=165 for the first one g(x)=150(1.03)^3=163.90

Vocaloid (vocaloid):

no, please look at f(x) and g(x) again...

OpenStudy (anonymous):

So f(x) is greater for the first 3 years am I wrong?

OpenStudy (anonymous):

oh sorry f(x)

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

I was talking to Nicole, sorry for the confusion

Vocaloid (vocaloid):

yes, f(x) is greater for 3 years now calculate f(x) and g(x) when x = 10

OpenStudy (anonymous):

g(x) is greater

OpenStudy (anonymous):

Yeah vocaloid would tell you to do f(x)=5(10)+150=200 g(x)=150(1.03)^10=201.5

OpenStudy (anonymous):

because i got 200 for f(x) and 201 for g(x)

OpenStudy (anonymous):

Yeah it's so close

Vocaloid (vocaloid):

right, so our final answer is...?

OpenStudy (anonymous):

a

Vocaloid (vocaloid):

yeah, good job ~

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Make sure you give a medal to vocalic

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

do you think you can help me with a couple more questions? @Vocaloid